I'm not sure what category to put this question in.
Solve the following equation for A: 2A/3 = 8 + 4A

Answer by stanbon(75887)

2A/3 = 8 + 4A
-----------------
(2/3)A = 8 + 4A
---
4A - (2/3)A = -8
---
(12/3)A - (2/3)A = -8
---
(10/3)A = -8
A = (3/10)(-8)
A = -12/5 = -2 2/5
=========================
Cheers,
Stan H.